# AutoV: Learning to Retrieve Visual Prompt for Large Vision-Language Models

### v1

@@ Abstract
Inspired by text prompts in large language models (LLMs), visual prompts have been explored to enhance the reasoning capabilities of large vision-language models (LVLMs). Current methods design heuristic visual prompts, such as overlaying a text-query-guided attention heatmap on the original input image. However, designing effective prompts manually is challenging and time-consuming, and it often fails to explore the benefits of different visual prompts, leading to sub-optimal performance. To this end, we propose AutoV that learns to automatically select the optimal visual prompt from various candidates based on given textual queries and the input image. To train AutoV, we developed an automatic data collection and labeling pipeline that evaluates various visual prompts with a pre-trained LVLM. We input a set of visual prompts into the LVLM and rank them according to the prediction losses generated by the model. Using the ranking as a supervision signal, we train AutoV to automatically choose the optimal visual prompt from various visual prompts for LVLMs. Experimental results indicate that AutoV enhances the performance of various LVLMs across multiple popular image understanding tasks. For instance, LLaVA-OV with AutoV achieves accuracy gain on LLaVA, and AutoV boosts Qwen2.5-VL by on MMMU, highlighting its potential as an optimal visual prompting method for LVLMs.
Shanghang Zhang at
@@ 1 Introduction
Recent advancements in large language models (LLMs) [39, 10, 1, 47, 37, 8] have spurred significant progress in large vision-language models (LVLMs), which connect visual encoders with language model decoders to effectively adapt textual reasoning capabilities to visual understanding tasks [31, 32, 4, 13, 26, 25, 5, 24]. Therefore, visual input plays a crucial role in LVLMs, enabling precise perception and fine-grained grounding across various multimodal scenarios.
@@
To further improve the perception and reasoning of LVLMs, visual prompting [55, 30, 53, 43] has emerged as an important technique analogous to textual prompting in LLMs [22, 65, 49]. Visual prompts enhance model performance by adding visual annotations such as masks, colorful circles, or heatmaps, thus guiding the attention of LVLMs toward crucial image regions of interest. However, existing visual prompting methods predominantly rely on heuristically designed prompts, limiting their ability to fully exploit the inherent diversity of visual prompts. As illustrated in Figure 1, we list the benchmark-level performance of existing heuristic prompt methods (e.g., FGVP [53], Red Circle [43], and API [55]). While all three methods demonstrate effectiveness on logic-intensive MMMU tasks, both blur-based FGVP and bounding prompts adversely affect vision-centric benchmarks (e.g., VizWiz [9]) and OCR-dependent tasks (e.g., TextVQA [44]). This is because heuristic prompt methods fail to automatically tune hyperparameters for each benchmark or even each instance.
@@
Hence, to develop a more effective visual prompt strategy suitable for diverse vision-language tasks, it is vital to adaptively retrieve the optimal visual prompt that most significantly improves the perception and reasoning of LVLMs in response to specific queries. Motivated by this, we introduce AutoV, an innovative visual prompting retrieval framework designed to dynamically select the optimal visual prompt from a diverse set of candidates, explicitly tailored to the textual query and input image. Specifically, we first generate a diverse set of visual prompt candidates111In this work, we use attention-based visual prompts from various transformer layers of CLIP in API [55]. and encode their representation. This diversity enables the framework to capture distinct visual representations that may be optimal for various query-image pairs. Next, we design a lightweight ranking network that efficiently integrates visual and textual information to predict the rankings preferences over prompt candidates. Finally, the highest-ranking candidate is chosen as the input for LVLMs during inference.
@@
To train the ranking network, we establish an automated pipeline for collecting and labeling training data, evaluating the rank of each visual prompt candidate based on the prediction losses of a pre-trained LVLM. Subsequently, the ranking network is supervised using a reward loss that differentiates between selected and rejected candidates. This strategy offers reliable supervision, allowing AutoV to attain adaptive, query-specific prompt retrieval without the need for extensive manual annotation.
@@
AutoV achieves significant performance improvements across various prominent LVLM architectures and diverse image understanding benchmarks, including fine-grained comprehension, real-world visual assistance, and complex open-ended reasoning tasks. For instance, integrating AutoV with LLaVA-OneVision and InternVL2 leads to improvements of and on LLaVA, respectively. Additionally, AutoV enhances Qwen2.5-VL on MMMU by 1.9%, highlighting its significant advantage in adaptive visual prompting retrieval. In summary, our contributions are as follows:
@@
We present AutoV, a novel automatic visual prompting framework for large vision-language models that adaptively retrieves the optimal visual prompt from a diverse candidate set, explicitly tailored to textual query and visual context pairs.
@@
We create a scalable and automated data collection pipeline equipped with a reward-based loss to effectively train a lightweight ranking network, facilitating accurate and adaptive visual prompt selection without the need for extensive manual annotations.
@@
We perform comprehensive experiments to validate the efficacy and generalizability of AutoV across various multimodal models and benchmarks, showcasing consistent and significant performance improvements over existing visual prompting methods. Once trained, AutoV integrates seamlessly into various LVLMs without additional fine-tuning.
@@ 2 Related Work
Large Vision-Language Models. Recent advances in large language models [39, 10, 1, 47, 37, 8] have catalyzed a new wave of large vision-language models [31, 2, 4, 13, 26, 25]. By coupling a high-capacity visual encoder with an LLM decoder, LVLMs transfer textual reasoning to visual understanding, and the quality of input images determines the lower bound of multimodal abilities. For instance, the popular LLaVA-OneVision [24] and Qwen2.5-VL [5] adopt a higher and dynamic resolution recipe for training to enhance their fine-grained reasoning, while vision-centric scaling in InternVL2 [11] enlarges the ViT [15] backbone to boost grounding capability. However, the above methods require training times of several hundred GPU days. Here, can we further boost the existing LVLMs by optimizing the visual input itself during the inference stage? The answer is Yes. Visual prompting [45, 29] is an emerging and effective technique in this direction.
@@
Visual Prompting for LVLMs. LVLMs relying solely on textual inputs exhibited two key limitations: visual hallucinations [6, 18] and language-driven biases [50, 38]. To better mitigate these challenges, visual prompting [60, 45, 29, 51] has emerged as an alternative way of supplying cues in the visual modality. Compared with textual prompts, it is more concise and explicit, while allowing guidance that reaches fine-grained, pixel-level details. Early work [43] shows that simply drawing a red circle around a target object steers CLIP [40] attention, yielding great performance on referring-expression and keypoint-localization benchmarks. FGVP [53] extends this idea with fine-grained prompting, replacing coarse shapes with blur-reversal prompts based on segmentation masks, further improving accuracy on RefCOCO [61]. AlphaCLIP [45] trains an augmented version of CLIP with an auxiliary alpha channel for indicating regions at the expense of increased training overhead. API [55] proposes attention prompting on images, overlaying text-guided attention heatmaps, boosting LVLMs on VQA, and open-ended reasoning tasks. Although previous studies have explored various visual prompts, it remains unclear which prompt is optimal for a specific image-question pair, leading to a decrease in the practicability of visual prompts. To address this limitation, we introduce our visual prompting method that learns to retrieve the optimal visual prompt for a specific image-question pair.
@@
Visual Retrieval for LVLMs. A variety of strategies have been proposed for visual retrievals, such as visual token selection [12, 63, 54] for LVLMs. Existing visual prompt selection pipelines generally adhere to three main paradigms. (1) The heuristic paradigm identifies benchmark-level optimal parameters through extensive evaluation but lacks instance-specific adaptation. (2) The random paradigm offers a stochastic baseline characterized by instance-level variation, but it lacks performance guarantees. This establishes it as the lower bound for adaptive approaches. (3) The Mixture-of-Experts (MoE) paradigm [16, 7, 66] provides the most direct form of adaptation by utilizing the GateNet specifically for visual retrieval. However, MoE gates are susceptible to overfitting during training, which can result in poor generalization. To mitigate these limitations, we propose an efficient ranking network specifically designed for visual retrieval, which models the pairwise preference ordering.
@@ 3 Proposed Approach: AutoV
In this section, we present AutoV for optimal visual prompt retrieval. Our framework includes four critical components: (1) feature extraction of prompt candidates, (2) candidate ranking consisting of modality interaction and single-modality mapping modules, (3) reward loss supervision along with automated data generation, and (4) an efficient inference pipeline. The overall architecture is illustrated in Figure 2, where the four modules are color-coded for clarity. Best viewed in color.
@@ 3.1 Feature Extraction of Prompt Candidates
The primary goal is to effectively choose the most appropriate visual prompt in response to various text queries and input images. For a group of input visual prompts , we employ a visual encoder (e.g., like CLIP [40]) to generate the visual feature:
@@ 3.2 Candidate Ranking Network
The goal of the ranking network is to effectively rank the visual prompt candidates based on their relevance to the textual query, by modeling the relationship between the visual tokens and the input text. Therefore, we first propose the modality interaction module to enable visual tokens to fuse the context information of text tokens. Inspired by observations in [62] that modality-fusion occurs in the early layers of the LLM, we use the first layer of the LLM decoder for modality interaction. Thus visual candidate tokens and text tokens are concatenated and fed into the interaction module:
@@
Here, only the vision and text mapping modules need additional training, each consisting of a single FFN with two linear layers. The total parameters are , resulting in lower training costs.
@@ 3.3 Reward Loss Supervision
Automated Data Generation. We create customized data to train the ranking network to select the optimal visual prompt for a given LVLM. Our approach is grounded in a simple intuition: a superior visual prompt should yield lower language modeling loss when input into any trained LVLM. We optimize the ranking network using relative loss instead of a one-hot selection vector, as correctness can stem from the LVLM’s language priors without relying on any input content [52, 56]. In contrast, the inference loss indicates the confidence of models in generating answers based on the input visual-question pair: a lower loss indicates higher confidence [17, 64].
@@
with an illustrative example provided in Figure 2, and the dataset we adopt is detailed in Appendix B.
Reward Loss Supervision. Inspired by the reward modeling from human feedback in OpenAI [36], we train our ranking network using reward modeling to align its partial order with the optimal visual prompt. Here, we model visual retrieval as a customized ranking task rather than a classification or regression problem. This is because the objective is not to assign absolute labels or scores, but to identify the most appropriate prompt from a set of candidates by comparing their relative effectiveness in enhancing multimodal understanding. Ranking allows the model to learn fine-grained preferences across diverse query-image combinations and better align with the real-world setting.
@@
3.4 Efficient Inference Pipeline
During the inference stage, we input several suitable visual prompt candidates along with the query into the large vision-language model and then pass them through our ranking network, as illustrated in Figure 2. Next, the ranking network outputs the corresponding reward scalar of each visual prompt. We then select the highest-scoring visual prompt, concatenate its visual tokens with text embeddings, and decode their representation through the LLM to produce the final answer.
@@ 4.1 Experiment Settings
Models. We apply our AutoV to four mainstream open-source large vision-language models, i.e., LLaVA-v1.5 [33], LLaVA-OneVision [24], Qwen2.5-VL [5], and InternVL2 [14]. In particular, LLaVA-v1.5 employs CLIP-pretrained ViT [40] as the visual tower and Llama2 [48] as the LLM backbone. LLaVA-OneVision enhances this design with SigLIP [59] for high-resolution refinement and replaces the LLM with Qwen2 [3]. Qwen2.5-VL features a dynamic resolution encoder, while InternVL2 adopts InternViT [13] for visual encoding and InternLM2 [11] as its LLM decoder.
Benchmarks. We evaluate AutoV on six widely-adopted image understanding benchmarks, including MMMU [58], POPE [27], VizWiz [9], MMVet [57], VQA (TextVQA) [44], and LLaVA (LLaVA-Bench in-the-wild) [31]. The evaluation benchmarks contain various aspects: hallucination suppression (POPE), STEM reasoning (MMMU), real-world visual assistance (VizWiz), fine-grained comprehension (MMVet), text-rich understanding (VQA), and open-world reasoning (LLaVA).

### v2

# Loss-Oriented Ranking for Automated Visual Prompting in LVLMs

@@ Abstract
Inspired by text prompts in large language models (LLMs), visual prompts have been explored to enhance the reasoning capabilities of large vision-language models (LVLMs). Current methods design heuristic visual prompts, such as overlaying a text-query-guided attention heatmap on the original input image. However, designing effective prompts manually is challenging and time-consuming, and it often fails to explore the benefits of different visual prompts, leading to sub-optimal performance. To this end, we propose AutoV that learns to automatically select the optimal visual prompt from various candidates based on given textual queries and the input image. To train AutoV, we develop an automatic data collection and labeling pipeline that evaluates various visual prompts with a pre-trained LVLM. We input a set of visual prompts into the LVLM and rank them according to the prediction losses generated by the model. Using the ranking as a supervision signal, we train AutoV to automatically choose the optimal visual prompt from various visual prompts for LVLMs. Experiments indicate that AutoV enhances the performance of various LVLMs across multiple image understanding tasks. For instance, LLaVA-OV with AutoV achieves accuracy gain on VizWiz, and AutoV boosts Qwen2.5-VL by on MMMU, highlighting its potential as an optimal visual prompting method.
@@ 1 Introduction
Recent advancements in large language models (LLMs) [brown2020language, achiam2023gpt, touvron2023llama, peng2023instruction, bi2024deepseek] have spurred significant progress in large vision-language models (LVLMs), which connect visual encoders with language model decoders to effectively adapt textual reasoning capabilities to visual understanding tasks [liu2023improvedllava, Qwen-VL, chen2023internvl, bai2025qwen2, li2024llava]. Therefore, visual input plays a crucial role in LVLMs, enabling precise perception and fine-grained grounding across various multimodal scenarios.
@@
To further improve the perception and reasoning of LVLMs, visual prompting [api, lin2024training, yang2023fine, shtedritski2023does] has emerged as an important technique analogous to textual prompting in LLMs [kojima2022large, zhou2022cocoop, wei2022chain]. Visual prompts enhance model performance by adding visual annotations such as masks, colorful circles, or heatmaps, thus guiding the attention of LVLMs toward crucial image regions of interest. However, existing visual prompting methods predominantly rely on heuristically designed prompts, limiting their ability to fully exploit the inherent diversity of visual prompts. As illustrated in Figure 1, we list the benchmark-level performance of existing heuristic prompt methods (e.g., FGVP [yang2023fine], Red Circle [shtedritski2023does], and API [api]). While all three methods demonstrate effectiveness on logic-intensive MMMU tasks, both blur-based FGVP and bounding prompts adversely affect vision-centric benchmarks (e.g., VizWiz [bigham2010vizwiz]) and OCR-dependent tasks (e.g., TextVQA [singh2019towards]). This is because heuristic prompt methods fail to automatically tune hyperparameters for each benchmark or even each instance.
@@
Hence, to develop a more effective visual prompt strategy suitable for diverse vision-language tasks, it is vital to adaptively retrieve the optimal visual prompt that most significantly improves the perception and reasoning of LVLMs in response to specific queries. Motivated by this, we introduce AutoV, an innovative visual prompting retrieval framework designed to dynamically select the optimal visual prompt from a set of candidates, explicitly tailored to the textual query and input image. Specifically, we first generate various visual prompt candidates and encode their representation. This diversity enables the framework to capture distinct visual representations that may be optimal for various query-image pairs. Next, we design a lightweight ranking network that efficiently integrates visual and textual information to predict the rankings preferences over prompt candidates. Finally, the highest-ranking candidate is chosen as the input for LVLMs during inference.
@@
To train the ranking network, we establish an automated pipeline for collecting and labeling training data, evaluating the rank of each visual prompt candidate based on the prediction losses of a pre-trained LVLM. Subsequently, the ranking network is supervised using a reward loss that differentiates between selected and rejected candidates. This strategy provides reliable supervision, allowing AutoV to attain adaptive, query-specific prompt retrieval without the need for extensive manual annotation work.
@@
AutoV achieves notable performance improvements across various prominent vision-language models and diverse image understanding benchmarks, including fine-grained comprehension, real-world visual assistance, and complex open-ended reasoning tasks. For instance, integrating AutoV with LLaVA-OneVision [li2024llava] and InternVL2 [chen2023internvl] leads to improvements of and on VizWiz [bigham2010vizwiz], respectively. Additionally, AutoV enhances Qwen2.5-VL [Qwen-VL] on MMMU [yue2024mmmu] by , highlighting its significant advantage in visual prompting reasoning retrieval. Furthermore, AutoV generalizes well, maintaining stable performance across captioning, grounding, and classification tasks. In summary, our contributions are as follows:
@@
We introduce AutoV, a novel automatic visual prompting framework for large vision-language models that adaptively retrieves the optimal visual prompt from a diverse candidate pool, explicitly tailored to each textual query.
@@
We create a scalable and totally automated data collection pipeline equipped with a reward-based loss to effectively train a lightweight ranking network, facilitating accurate and adaptive visual prompt selection without the need for extensive manual annotations.
@@
We perform comprehensive experiments to validate the efficacy and robustness of AutoV across various multimodal models and tasks, showcasing consistent and significant improvements over existing visual prompting methods. Once trained, AutoV integrates seamlessly into various LVLMs without additional fine-tuning.
@@ 2 Related Work
Large Vision-Language Models. Recent advances in large language models [radford2019language, brown2020language, achiam2023gpt, touvron2023llama, peng2023instruction, bi2024deepseek] have catalyzed a new wave of large vision-language models [liu2023improvedllava, alayrac2022flamingo, Qwen-VL, chen2023internvl, li2024mini, li2023blip]. By coupling a high-capacity visual encoder with an LLM decoder, LVLMs transfer textual reasoning to visual understanding, and the quality of input images determines the lower bound of multimodal abilities. For instance, the popular LLaVA-OneVision [li2024llava] and Qwen2.5-VL [bai2025qwen2] adopt a higher and dynamic resolution recipe for training to enhance their fine-grained reasoning, while vision-centric scaling in InternVL2 [cai2024internlm2] enlarges the ViT [dosovitskiy2020image] backbone to boost grounding capability. However, the above methods require training times of several hundred GPU days. Here, can we further boost the existing LVLMs by optimizing the visual input itself during the inference stage? The answer is Yes. Visual prompting [sun2023alphaclip, lin2024drawandunderstand, jiang2024joint, zhang2024exploring, api, yang2023fine, shtedritski2023does] is an emerging and effective technique in this direction.
@@
Visual Prompting for LVLMs. LVLMs relying solely on textual inputs exhibited two key limitations: visual hallucinations [bai2024hallucination, huang2024visual] and language-driven biases [wu2024commit, qu2024unified]. To better mitigate these challenges, visual prompting [zhang2023vpgtrans, sun2023alphaclip, lin2024drawandunderstand, wu2024visual] has emerged as an alternative way of supplying cues in the visual modality. Compared with textual prompts, it is more concise and explicit, while allowing guidance that reaches fine-grained, pixel-level details. Early work [shtedritski2023does] shows that simply drawing a red circle around a target object steers CLIP [CLIP] attention, yielding great performance on referring-expression and keypoint-localization benchmarks. FGVP [yang2023fine] extends this idea with fine-grained prompting, replacing coarse shapes with blur-reversal prompts based on segmentation masks, further improving accuracy on RefCOCO [zhang2019referring]. AlphaCLIP [sun2023alphaclip] trains an augmented version of CLIP with an auxiliary alpha channel for indicating regions at the expense of increased training overhead. API [api] proposes attention prompting on images, overlaying text-guided attention heatmaps, boosting LVLMs on VQA, and open-ended reasoning tasks. Although previous studies have explored various visual prompts, it remains unclear which prompt is optimal for a specific image-question pair, leading to a decrease in the practicability of visual prompts. To address this limitation, we introduce our visual prompting method that learns to retrieve the optimal visual prompt for a specific image-question pair.
@@
Visual Retrieval for LVLMs. A variety of strategies have been proposed for visual retrievals, such as visual token selection [chen2024image, zhang2024sparsevlm, yang2024visionzip] for LVLMs. Existing visual prompt selection pipelines generally adhere to three main paradigms. (1) The heuristic paradigm identifies benchmark-level optimal parameters through extensive evaluation without instance-specific adaptation. (2) The random paradigm offers a baseline characterized by instance-level variation. (3) The Mixture-of-Experts (MoE) paradigm [fedus2022switch, bao2022vlmo, zong2024mova] provides the most direct form of adaptation by utilizing the GateNet specifically for visual retrieval, which is susceptible to overfitting during training.
Notably, our AutoV fundamentally differs from retrieval-augmented methods [qi2024rora, xing2025re]. Unlike them, which retrieve supplementary visual data from external databases and require model finetuning, AutoV selects the most suitable visual prompt based solely on the internal features of the current image through a lightweight ranking network. Moreover, it operates within a compact prompt pool rather than a large-scale database, yielding substantial efficiency gains.
@@ 3 Proposed Approach: AutoV
In this section, we propose the AutoV for optimal visual prompt retrieval. Our framework includes four critical components: (1) feature extraction of prompt candidates, (2) candidate ranking consisting of modality interaction and single-modality mapping modules, (3) reward loss supervision along with automated data generation pipeline, and (4) an efficient and robust inference pipeline. The overall architecture is illustrated in Figure 2, where the four modules are color-coded for clarity. Best viewed in color.
@@ 3.1 Feature Extraction of Prompt Candidates
The primary goal is to effectively choose the most appropriate visual prompt in response to various text queries and input images. For a group of input visual prompts , we employ a visual encoder (e.g., like CLIP [CLIP]) to generate the visual feature:
@@ 3.2 Candidate Ranking Network
The goal of the ranking network is to effectively rank the visual prompt candidates based on their relevance to the textual query, by modeling the relationship between the visual tokens and the input text. Therefore, we first propose the modality interaction module to enable visual tokens to fuse the context information of text tokens. Inspired by observations in [zhang2025llava] that modality-fusion occurs in the early layers of the LLM, we use the first layer of the LLM decoder for modality interaction. Thus, visual candidate tokens and text tokens are concatenated and fed into the interaction module:
@@
where denotes the th visual prompt candidate. Here, only the vision and text mapping modules need additional training, each consisting of a single FFN with two linear layers. The modules comprise parameters in total, thereby resulting in lower training costs.
@@ 3.3 Reward Loss Supervision
Automated Data Generation. We create customized data to train the ranking network to select the optimal visual prompt for a given LVLM. Our approach is grounded in a simple intuition: a superior visual prompt should yield lower language modeling loss when input into any trained LVLM. We optimize the ranking network using relative loss instead of a one-hot selection vector, as correctness can stem from the LVLM’s language priors without relying on any input content [xiao2024can, yu2024frame]. In contrast, the inference loss indicates the confidence of models in generating answers based on the input visual-question pair: a lower loss indicates higher confidence [guo2017calibration, zhang2024unveiling].
@@
with an illustrative example provided in Figure 2, and the dataset we adopt is detailed in Appendix A.
Reward Loss Supervision. Inspired by the reward modeling from human feedback in OpenAI [ouyang2022training], we train our ranking network using reward modeling to align its partial order with the optimal visual prompt. Here, we model visual retrieval as a customized ranking task rather than a classification or regression problem. This is because the objective is not to assign absolute labels or scores, but to identify the most appropriate prompt from a set of candidates by comparing their relative effectiveness in enhancing multimodal understanding. Ranking allows the model to learn fine-grained preferences across diverse query-image combinations and better align with the real-world setting.
@@
3.4 Robust Inference Pipeline
During the inference, we input multiple candidate visual prompts along with the query into the large vision-language model and pass them through our ranking network, as shown in Figure 2. The network outputs a reward scalar for each visual prompt, and we select the highest-scoring one. Its visual tokens are concatenated with the text embeddings and decoded by the LLM to generate the final answer.
Besides, to reduce distributional bias, we introduce a pre-filtering step: the prompt whose visual features are farthest from other candidates, as measured by cosine distance, is removed in advance, further mitigating the negative effects of suboptimal prompts and enhancing robustness.
@@ 4.1 Experiment Settings
Models. We apply our AutoV to four mainstream open-source large vision-language models, i.e., LLaVA-1.5 series [liu2024visual], LLaVA-OneVision [li2024llava], Qwen2.5-VL [bai2025qwen2], and InternVL2 [chen2024expanding]. In particular, LLaVA-1.5 series build on CLIP-ViT [CLIP] and Llama2 [touvron2023llama2]. LLaVA-OneVision enhances it with SigLIP [siglip] and Qwen2 [bai2023qwen]. Qwen2.5-VL introduces a dynamic-resolution encoder, and InternVL2 integrates InternViT [chen2023internvl] with InternLM2 [cai2024internlm2].
Benchmarks. AutoV is extensively evaluated on fifteen vision-language benchmarks, covering video reasoning (CVBench [tong2024cambrian]), vision-centric tests (BLINK [fu2024blink], MMVP [tong2024eyes], MMStar [chen2024we]), comprehensive multimodal evaluation (MM-Vet [yu2023mm]), knowledge/math reasoning (MMMU [yue2024mmmu], MathVista [lu2023mathvista]), robustness and VQA (VizWiz [bigham2010vizwiz], Real-World QA [RealworldQA], TextVQA [singh2019towards], LLaVA-Wild [liu2023improvedllava]), grounding (RefCOCO+ [yu2016modeling]), object-hallucination probing (POPE [li2023evaluating]), captioning (COCO-Caption2017 [lin2014microsoft]), and multimodal classification (ImageNet-1K [deng2009imagenet]).